# SANA-Video: Efficient Video Generation with Block Linear Diffusion Transformer

### v1

@@ 1 Introduction
Block Linear Attention with KV Cache. The success of SANA-Video in long video generation is mainly inspired by the attribute of causal linear attention [11]. Based on our reformulation of the causal linear attention operation, we reduce the KV cache to a small and fixed memory, along with a fixed computational cost for each new token. This natively supports long-context operations. Based on the block linear attention module, we introduce a two-stage autoregressive model continue-training paradigm, including autoregressive block training with monotonically increasing SNR sampler and the improved self-forcing specially for our long context attention operation, leading to efficient, long, and high quality video generation.
@@ 3.2 Efficient Linear DiT Pre-Training
3.3 Block Linear Attention
@@ 3.3.1 Block Linear Attention with KV Cache
Block Causal Mix-FFN. In addition to linear attention, In addition to linear attention, our designed temporal-spatial Mix-FFN with convolution layers improves locality and must also be causal for long video generation. During inference, the causal temporal convolution layer with a kernel size of 3 requires access to the last frame of the previous block. To this end, the last frame of previous block is also cached (Fig. 4(b)). Therefore, our causal linear DiT module keeps a fixed memory cache, containing cumulative sum of attention states and keys from all previous frames, along with the last frame of the previous block.
@@ 3.3.2 Autoregressive Block Training
The continue training of the autoregressive SANA-Video variant begins with the pre-trained 5s SANA-Video model. To align with the pre-trained model’s distribution, we propose a monotonically increasing SNR sampler. Specifically, we randomly select a block and sample a timestep for it with the SNR sampler [14]. Then the timesteps for the remaining blocks are sampled via propagated probability [19], ensuring all the timesteps are monotonically increasing, i.e., later blocks have larger timestep than early blocks. This proposed timestep sampler offers two key advantages. First, the monotonically increasing timesteps have a much smaller sampling space than random timesteps, which results in faster convergence and better performance. Second, applying the SNR sampler to a randomly selected block guarantees that every block is trained with sufficient information.
However, monotonically increasing SNR sampler cannot address a severe problem in autoregressive generation, i.e., exposure bias, where condition blocks are ground truth during training but are generated content during inference, leading to error accumulation and limiting performance in long video generation. Self-Forcing [17] aims to address this issue in a vanilla attention DiT model with autoregressive rollout. In our work, we improve Self-Forcing to better leverage our constant-memory global KV cache. Limited by the increasing VRAM requirement of causal vanilla attention (Fig. 1(c) and Table 1), Self-Forcing uses local attention within a designed window size. Consequently, it sets the length of self-generated content to be the same as the pre-trained model (i.e., 5s). In contrast, the block linear attention in SANA-Video supports a long-context global KV cache with a small and constant GPU memory. This allows us to self-generate a much longer video (e.g., 1 min) and select a clip for training, which better aligns the conditioning signals between training and inference.
3.3.3 Block Linear Attention during Inference
We follows Self-Forcing [17] for autoregressive inference, with the KV cache update based on our design (Alg. 1). Specifically, we first initialize KV cache as empty and start to denoise the first block. After it is fully denoised, the attention state , cumulative sum of keys and cache for convolustion in Spatial-Temporal Mix-FFN (conv cache ) will be stored. For the remaining blocks (e.g., -th block), they will use the existing KV cache to denoise the latent until clean and then update the cumulative attention state and cumulative sum of keys . Also, conv cache will be replaced with the new cache. Such update leverages the global while keeping the memory constant and small, making the long video generation efficient and effective.
@@ C.4 Influence of Motion Score
C.5 Long Video Visualization
In Fig. 12, we provide an example of our 30-second long video generation. SANA-Video is able to generate motion consistent and semantically aligned long videos.

### v2

@@ 1 Introduction
Block Linear Attention with KV Cache. The success of SANA-Video in long video generation, a.k.a, LongSANA, is mainly inspired by the attribute of causal linear attention [11]. Based on our reformulation of the causal linear attention operation, we reduce the KV cache to a small and fixed memory, along with a fixed computational cost for each new token. This natively supports long-context operations. Based on the block linear attention module, we introduce a two-stage autoregressive model continue-training paradigm, including autoregressive block training with monotonically increasing SNR sampler and the improved self-forcing specially for our long context attention operation, leading to efficient, long, and high quality video generation.
@@ 3.2 Efficient Linear DiT Pre-Training
3.3 LongSANA with Block Linear Attention
@@ 3.3.1 Block Linear Attention with KV Cache
Block Causal Mix-FFN. In addition to linear attention, our proposed temporal-spatial Mix-FFN enhances locality using convolutional layers. To support long video generation, this module must also operate causally. We ensure causal processing during both training and inference with two operations, as illustrated in Fig. 4(b). First, to prevent information leakage from subsequent blocks during training, we append an all-zero token (‘Zero Padding’ ) to the end of each block . Second, our causal temporal convolution (kernel size 3) requires the last frame of the preceding block. We address this by caching the last token of each block (‘’ ) and prepending it to the next. Overall, our causal linear DiT module keeps a fixed memory cache, containing cumulative sum of attention states and keys from all previous frames for attention, along with the last frame of the previous block for Mix-FFN.
@@ 3.3.2 Autoregressive Block Training
The continue training of the autoregressive SANA-Video variant, i.e. LongSANA, begins with the pre-trained 5s SANA-Video model. To align with the pre-trained model’s distribution, we propose a monotonically increasing SNR sampler. Specifically, we randomly select a block and sample a timestep for it with the SNR sampler [14]. Then the timesteps for the remaining blocks are sampled via propagated probability [19], ensuring all the timesteps are monotonically increasing, i.e., later blocks have larger timestep than early blocks. This proposed timestep sampler offers two key advantages. First, the monotonically increasing timesteps have a much smaller sampling space than random timesteps, which results in faster convergence and better performance. Second, applying the SNR sampler to a randomly selected block guarantees that every block is trained with sufficient information.
However, monotonically increasing SNR sampler cannot address a severe problem in autoregressive generation, i.e., exposure bias, where condition blocks are ground truth during training but are generated content during inference, leading to error accumulation and limiting performance in long video generation. Self-Forcing [17] aims to address this issue in a vanilla attention DiT model with autoregressive rollout. Limited by the increasing VRAM requirement of causal vanilla attention (Fig. 1(c) and Table 1), Self-Forcing uses local attention within a designed window size. Consequently, it sets the length of self-generated content to be the same as the pre-trained model (i.e., 5s). Later on, LongLive [20] explores streaming long training on 1 minute video, but it still limits to the local attention with sink due to the complexity of full attention. In contrast to the full attention, the block linear attention in LongSANA supports a long-context global KV cache with a small and constant GPU memory. This allows us to further extend LongLive with global attention when self-generating a much longer video (e.g., 1 min), which better aligns the conditioning signals between training and inference and keeps better temporal consistency.
3.3.3 Real-time Long Video Generation with Block Linear Attention
We follows Self-Forcing [17] for autoregressive inference, with the KV cache update based on our design (Alg. 1). Specifically, we first initialize KV cache as empty and start to denoise the first block. After it is fully denoised, the attention state , cumulative sum of keys and cache for convolustion in Spatial-Temporal Mix-FFN (conv cache ) will be stored. For the remaining blocks (e.g., -th block), they will use the existing KV cache to denoise the latent until clean and then update the cumulative attention state and cumulative sum of keys . Also, conv cache will be replaced with the new cache. Such update leverages the global while keeping the memory constant and small, making the long video generation efficient and effective. Attribute to the efficient block linear attention, our 4-step LongSANA is able to generate 1-min and 16 FPS 480P video within 35 seconds on NVIDIA H100 GPU, achieving real-time, 27 FPS generation speed.
@@ C.4 Influence of Motion Score
C.5 LongSANA Visualization
In Fig. 12, we provide an example of our 1-minute long video generation. LongSANA is able to generate motion consistent and semantically aligned long videos.